OpenStudy (anonymous):

Find g'(4) given that f(4)=3 and f'(4)=-5 g(x)=sqrt(x) * f(x) Use producst rule

myininaya (myininaya):

use the product rule

myininaya (myininaya):

you know the derivative of sqrt(x) and the derivative of f(x) right?

myininaya (myininaya):

derivative of f(x) is f'(x) derivative of sqrt(x) is ? recall the product rule is \[(u(x) \cdot v(x))'=u'(x) \cdot v(x)+u(x) \cdot v'(x)\]

OpenStudy (anonymous):

yes I know that the derivative of sqrt(x) is 1/2sqrt(x). I get (-5x^1/2+3)/2sqrt(x) . but that is not the right answer